The owner of Genuine Subs, Inc., hopes to expand the present operation by adding one new outlet. She has studied three locations. Each would have the same labor and materials costs (food, serving containers, napkins, etc.) of $1.80 per sandwich. Sandwiches sell for $2.60 each in all locations. Rent and equipment costs would be $5,350 per month for location A, $5,675 per month for location B, and $5,925 per month for location C.

a. Determine the volume necessary at each location to realize a monthly profit of $9,750. (Do not round intermediate calculations. Round your answer to the nearest whole number.)

Location Monthly Volume
A
B
C


b-1. If expected sales at A, B, and C are 20,750 per month, 22,750 per month, and 23,750 per month, respectively, calculate the profit of the each locations? (Omit the "$" sign in your response.)

Location Monthly Profits
A $
B $
C $


b-2. Which location would yield the greatest profits?

  • Location B

  • Location A

  • Location C
Answer #1

a) Location A = Fixed cost + expected profits)/ (Selling price – Cost price)

= (5350 + 9750)/ (2.6-1.8)

= 18875 units

Location B = Fixed cost + expected profits)/ (Selling price – Cost price)

= (5675 + 9750)/ (2.6-1.8)

= 19281.25 = 19281 units

Location C = Fixed cost + expected profits)/ (Selling price – Cost price)

= (5925 + 9750)/ (2.6-1.8)

= 19593.75 = 19594 units

b1) Sales at A = 20750

Selling cost for sold units = (Sales/Selling price) *Cost Price

(20750/2.6) *1.8

14365.38 $

Profit for location A = Sales – cost of goods sold – Fixed costs

=20750 - 14365.38 - 5350

= 1034.62 $

Sales at B = 22750

Selling cost for sold units = (Sales/Selling price) *Cost Price

(22750/2.6) *1.8

15750 $

Profit for location B = Sales – cost of goods sold – Fixed costs

=22750 - 15750 - 5675

= 1325 $

Sales at C = 20750

Selling cost for sold units = (Sales/Selling price) *Cost Price

(23750/2.6) *1.8

16442.30 $

Profit for location C = Sales – cost of goods sold – Fixed costs

=23750 – 16442.30 – 5925

= 1382.7 $

b2) Location C will yield greates profits.